OpenStudy (anonymous):

factor x^2 -64?? PLEASEEEEEEEEEEEE

OpenStudy (anonymous):

x^2 = 64 hence x= 8 or -8

OpenStudy (callisto):

Use difference of 2 squares x^2 -64 = x^2 - 8^2 = (x-8)(x+8)

OpenStudy (diyadiya):

this is of the form a^2-b^2 a^2-b^2=(a-b)(a+b) x^2-64=x^2-8^2= (x-8)(x+8)